Question Help The pulse rates of 165 randomly selected adult males vary from a low of 43 bpm to a high of 107 bpm. Find the minimum sample size required to estimate the mean pu se rate of adult males. A sume that we want 9 % con dence that the sa k ra is within 2 bpm of the populat on mean. Complete parts a.Find the sample size using the range rule of thumb to estimate σ (Type a whole number) F3 F5 F10 F11 F12 Prtsk Pause
Answer #1

Given :-

Low = 43 bpm and High = 107 bpm

range = High - Low = 107 - 43 = 64

Therefore ,

Standard deviation ( sigma ) =( range/4 )= (64/4) = 16

90% confidence level at alpha =10% then value of z is 1.645

And margin of error (ME) = 2 bpm

Now to find the sample size.

ME = z * sigma /√n

=> n = [ z*sigma / ME ]2

=> n = [ 1.645*16 / 2 ]2

n = 174